Provide the metal oxidation state and dn electron configuration for the following complexes (Suggestions: add any lone pairs missing from the ligands before continuing!)

a) [Co(NH3)5(NO2)]2+

b) [ReH9]2-

c) [CuCl5]3-

d) [Cr(SCN)4(NH3)2]-
Answer #1

the metal oxidation state and dn electron configuration for the given complexes are as follows:

a) [Co(NH3)5(NO2)]2+

let assume the oxidation number of Co is X.

X+0+ (-1) =2+

X=3+

Electron configuration Co: [Ar] 3d7 4s2

Electron configuration Co3+ : [Ar] 3d6 4s0

b) [ReH9]2-

let assume the oxidation number of Re is X.

X+9(-1)=2-

X=7+

Electron configuration Re: [Xe] 6s2 4f14 5d5

Electron configuration Re7+ : [Xe] 6s0 4f14 5d0

c) [CuCl5]3-

let assume the oxidation number of Cu is X.

X+5(-1) =3-

X=2+

Electron configuration Cu: [Ar] 3d10 4s1

Electron configuration Cu2+ : [Ar] 3d9 4s0

d) [Cr(SCN)4(NH3)2]-

let assume the oxidation number of Cr is X.

X+4(-1)+ 0 =1-

X=3+

Electron configuration Cr: [Ar] 3d5 4s1

Electron configuration Cr3+ : [Ar] 3d3 4s0